OpenStudy (anonymous):

1 over 3 (2x - 8) = 4 solve for X can it also be explained?

OpenStudy (anonymous):

i got 6

OpenStudy (anonymous):

just need to double check

OpenStudy (mokeira):

\[\frac{ 1 }{3(2x-8) }=4\] First step is to remove the denominator by multiplying both sides by 3(2x-8) \[1=4 \times 3(2x-8)\] \[1=4(6x-24)\] \[1=24x-96\] \[1+96=24x\] \[97=24x\] do you have choices?

OpenStudy (anonymous):

yes 2, 6, 10, and 10 over 3

OpenStudy (mokeira):

none of the above but 10 over 3 is close

OpenStudy (anonymous):

oh ok thank you again!